# Test-Time Efficient Pretrained Model Portfolios for Time Series Forecasting

### v1

@@ G.4 Ablation Study
Portfolio Type freq 0.009 0.015 0.007 0.046
@@
Model Size 2m 0.017 0.014 0.018 0.017
Model Size 4m 0.018 0.007 0.018 0.008
Model Size tiny 0.016 0.001 0.014 0.006
@@
Portfolio Type domain -0.008 0.015 -0.005 0.088
Portfolio Type freq 0.007 0.032 0.008 0.002
Model Size 1m -0.002 0.695 -0.003 0.554
Model Size 2m -0.009 0.043 0.004 0.338
Model Size 4m -0.001 0.779 0.009 0.055
Model Size tiny 0.010 0.049 -0.004 0.224
Overall N/A -0.001 0.791 0.002 0.533
@@ G.5 Detailed Benchmark Results
While we provide the aggregated results in the main text, in this section, we provide detailed results for each individual dataset included in the benchmarks. Specifically, for BM2, we report the results for WQL and MASE in Tables 9 and 10, respectively. For GIFT-Eval, WQL results are shown in Tables 11 and 12, while MASE results are presented in Tables 13 and 14.

### v2

@@ G.4 Ablation Study
Portfolio Type freq 0.009 0.015 0.007 0.0536
@@
Model Size 2m 0.017 0.015 0.018 0.0238
Model Size 4m 0.018 0.01225 0.018 0.014
Model Size tiny 0.016 0.0023 0.014 0.014
@@
Portfolio Type domain -0.008 0.08575 -0.005 0.20533
Portfolio Type freq 0.007 0.08575 0.008 0.014
Model Size 1m -0.002 0.791 -0.003 0.554
Model Size 2m -0.009 0.08575 0.004 0.4732
Model Size 4m -0.001 0.791 0.009 0.1925
Model Size tiny 0.010 0.08575 -0.004 0.392
Overall N/A -0.001 0.791 0.002 0.554
G.4.1 Portfolios trained from scratch
In the result Table 2 of our ablation study, we show that there is no statistically significant difference in accuracy between portfolios trained from scratch and portfolios trained with post-training. In Table 9, we support this result by a rework of Table 1 that also includes portfolios of models that are trained from scratch, as opposed to post-training.
Table 9: Aggregated relative WQL and MASE on BM2 including portfolios trained from scratch, scaled against a single generalist model of size 1m.
WQL MASE
1m 2m 4m tiny 1m 2m 4m tiny
Model Selection 0.963 0.950 0.939 0.923 0.973 0.948 0.933 0.930
Model Selection 0.918 0.916 0.880 0.909 0.920 0.899 0.887 0.893
Domain (scratch) Ensemble (Greedy) 0.953 0.927 0.923 0.911 0.947 0.924 0.925 0.905
Model Selection 0.987 0.928 0.943 0.919 0.968 0.936 0.942 0.919
Frequency (scratch) Ensemble (Greedy) 0.907 0.905 0.893 0.910 0.906 0.895 0.882 0.880
Model Selection 0.935 0.914 0.885 0.940 0.920 0.913 0.895 0.896
Model Selection 0.990 0.966 0.942 0.944 0.979 0.961 0.935 0.941
It can be seen that there is no significant difference between these two type of portfolios in terms of accuracy. Similarly, from the individual tasks point of view, we also were not able draw a direct conclusion about how from-scratch and post-trained models compare. Despite this performance similarity, the portfolios trained from scratch required 200x more compute for each specialist.
@@ G.5 Detailed Benchmark Results
While we provide the aggregated results in the main text, in this section, we provide detailed results for each individual dataset included in the benchmarks. Specifically, for BM2, we report the results for WQL and MASE in Tables 10 and 11, respectively. For GIFT-Eval, WQL results are shown in Tables 12 and 13, while MASE results are presented in Tables 14 and 15.
@@
Appendix H Use of LLMs
Large language models are only used for a few word choices in the text. They are not used for the implementation of this research.